• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Roche

    Log entry time 00:00:39 on June30,2001

    Entry number 65039

    keyword=Shift Summary

    17:55 -> beam is back :
    1/2wp beam on (starting now)
    beam position ok, raster on, fast feed back working
    target status :
    field dir = 180deg
    laser conf long
    1/2 wp in
    grad coil l=-2.00A, T=5.00 A

    17:56->EPR measurement (37.5%)

    18:00 -> 21:00 production, the beam charge assymetry is large and we don't succeed to bring it low. After a while we determine that the rotatable waveplate seems not to move (read back stucked at 599). MCC will call the injector expert.

    21:00 -> Joe Grames comes and fix the trouble with the feedback on the charge assymetry (see entry # 65023)

    runs taken during this shift : 847 -> 854

    status : number of electrons at the end of runs 853
    21969 on left arm
    20107 on right arm